OpenStudy (anonymous):

help please! [9.01] What is the value of the coefficient "b" when the quadratic equation y = (3x + 2)(2x − 1)is written in standard form? please explain −7 −2 1 6

myininaya (myininaya):

so multiply (3x+2)(2x-1) and i bet you are looking for the coefficient of x

OpenStudy (anonymous):

i says coefficient of b?

myininaya (myininaya):

there is no b... ax^2+bx+c I think they are calling b the coefficient of x

myininaya (myininaya):

you need to expand (3x+2)(2x-1) by multiplying

OpenStudy (anonymous):

how do i do that?

myininaya (myininaya):

multiply?

myininaya (myininaya):

have you heard of the distributive property? (a+b)(c+d) =a(c+d)+b(c+d)

OpenStudy (anonymous):

how do i multipy them together? because theres two x's

myininaya (myininaya):

distribute and add like terms

OpenStudy (anonymous):

im so confused

myininaya (myininaya):

example \[(ax+b)(cx+d) \\ =ax(cx+d)+b(cx+d) \\ =ax \cdot cx + ax \cdot d+b \cdot cx +b \cdot d \\ =ac x^2+ax d+bc x+bd \\ =acx^2+(ad+bc)x+bd\\ \text{ the coefficent } \text{ of x here is } ad+bc\]

OpenStudy (anonymous):

so... (3x+2)(2x+1)?

myininaya (myininaya):

you need to first multiply

myininaya (myininaya):

have you done that yet

myininaya (myininaya):

you can follow my example if you want

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

multiply what

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

(3x+2)(2x+1) follow what I said... the first line of what I said is 3x(2x+1)+2(2x+1) now use the distributive property on both like I did in my third line

OpenStudy (anonymous):

im still lost

OpenStudy (anonymous):

@myininaya use more simply terms you use such big words

myininaya (myininaya):

distributive property?

myininaya (myininaya):

this is the name of the property

myininaya (myininaya):

a(b+c)=ab+ac

OpenStudy (anonymous):

i got it. its 1

myininaya (myininaya):

great